OpenStudy (anonymous):

Which of the following is a solution for the system y = -3x + 5 and y = 2x + 5? A. (1, 7) B. (5, 0) C. (2, -1) D. (0, 5)

OpenStudy (anonymous):

(d)

OpenStudy (callisto):

y = -3x + 5 -(1) y = 2x + 5 -(2) (2) - (1) 0 = 2x + 5 - ( -3x + 5 ) 0 = 5x x = 0 y = 2x +5 = 2(0) + 5 = 5 So the answer is (0, 5)